# Evaluating A/B Testing Methodologies via Sample Splitting: Theory and Practice

### v1

@@ 1 Introduction
Measuring the impact of proposed improvements to the way in which A/B tests are analyzed and used to inform decision making is critical for large-scale A/B testing products. Over the last few decades, there have been several proposed improvements to A/B testing methodologies, including new estimators of impacts (for example, [deng_etal_2013], [deng_etal_2023], and [deng_etal_2024]) and new rules for deciding when to launch (for example, [goldberg_johndrow_2017], [azevedo_etal_2019] and [azevedo_etal_2020]). The scope of proposed improvements continues to expand, with more recent work exploring AI-driven approaches (for example, [delriochanona_etal_2023] and [delriochanona_etal_2025]). A/B testing products must determine which of these methodologies deliver meaningful value relative to their implementation costs.
@@
Sample splitting offers a potential solution to this problem ([tripuraneni_etal_2023]), but its theoretical foundation remains undeveloped. Under sample splitting, each A/B test is divided into two, non-overlapping splits: one to implement the proposed methodology, the other to construct an independent, unbiased proxy for the true impact. Comparing these two quantities across A/B tests allows us to measure the average performance of the proposed methodology, without making strong distributional assumptions ([tripuraneni_etal_2023]). However, the statistical properties of these sample-split estimators have yet to be characterized formally: their bias, variance, and asymptotic distributions remain unknown, and there are no established methods for conducting statistical inference. Without such foundations, products lack guidance on exactly how to use sample splitting, limiting its adoption in practice.
@@
More generally, our work contributes to the literature on cross-validation and resampling methods ([stone_1971]). Like cross-validation, sample splitting separates the application of a methodology from its evaluation by partitioning data. The key difference is that in cross-validation the prediction target is observed directly, whereas in our setting the prediction target is itself an estimate subject to sampling variation. To our knowledge, [tripuraneni_etal_2023] were the first to apply the idea of sample splitting in the A/B testing context, building upon earlier work in the estimation of heterogeneous treatment effects by [athey_imbens_2016] and [chernozhukov_2025]. [tripuraneni_etal_2023] demonstrate the potential of sample splitting in evaluating the performance of A/B testing methodologies but leave its statistical properties largely unformalized. [^1] The authors acknowledge that their work is guided primarily by intuition and heuristics rather than principled theory. For example, they note the use “$t$-test heuristics” (page 7) and suggest that the confidence intervals they report “should be interpreted cautiously” (page 11). We extend their work by introducing a theoretical framework for sample splitting in A/B testing environments, characterizing both what these methods can measure and how to conduct valid inference.
@@ 2.1 Context
We consider a set of $i\in\{1,...,I\}$ feature changes, each evaluated through an A/B test where the change is exposed to a random subset of units. Each feature change has a true impact $\Delta_{i}$, which is a random variable with realized values denoted by $\delta_{i}$. These true impacts $\Delta_{i}$ are independently and identically distributed (iid) according to some unknown probability distribution $G$. Conditional on the true impact $\Delta_{i}=\delta_{i}$, the A/B test generates random data $\hat{X}_{i}$ independently distributed according to a probability distribution $F$, with realized values $\hat{x}_{i}$.
We define an A/B testing methodology $\kappa(\hat{x}_{i})$ to be a mapping of the data $\hat{x}_{i}$ to a real-valued output. The methodology could be an estimator of the true impact $\delta_{i}$ or a decision rule that determines whether to launch the given treatment ($\kappa(\hat{x}_{i})=1$) or not ($\kappa(\hat{x}_{i})=0$). Given the randomization of the feature change and the central limit theorem, we have an unbiased estimator $\kappa(\hat{X}_{i})=\hat{\Delta}_{i}$ of $\delta_{i}$ that is normally distributed, with realized values $\hat{\delta}_{i}$ and sampling variance $\tau_{i}^{2}$:
While the sampling variances $\tau_{i}^{2}$ are typically estimated from data, we follow standard practice in the A/B testing literature and treat them as known (see, for example, [azevedo_etal_2020]). In most large-scale A/B testing environments, $\tau_{i}^{2}$ is estimated with far greater precision than $\hat{\Delta}_{i}$, making this a pragmatic simplification that helps focus attention on the first-order uncertainty in $\hat{\Delta}_{i}$. [^2] In an A/B test with normally distributed outcomes and a total of $N$ units divided equally across treatment and control, the sampling variance of $\hat{\Delta}_{i}$ decreases at a rate of $1/N$, while the sampling variances of standard estimators of $\tau_{i}^{2}$ decrease at a rate of $1/N^{2}$. So, for example, when $N=1M$ the sampling variance of $\hat{\Delta}_{i}$ will be at least 1M times larger than the sampling variance of standard estimators of $\tau_{i}^{2}$. Similar results hold for heavy-tailed outcomes when winsorization is applied, as is common in large-scale A/B testing.
@@
In what follows, for any random variable $K_{i}$, we write $F_{K}$ to denote its distribution induced by the sampling distribution $F$ and $G_{K}$ to denote its distribution induced by the distribution of true impacts $G$. We write $E_{F}(\cdot)$ and $\text{Var}_{F}(\cdot)$ for the expectation and variance with respect to randomness induced by $F$, and $E_{G}(\cdot)$ and $\text{Var}_{G}(\cdot)$ for the expectation and variance with respect to randomness induced by $G$. Finally, we write $E(\cdot)$ and $\text{Var}(\cdot)$ for the expectation and variance with respect to both sources of randomness, integrating over both $F$ and $G$. We denote by $\phi(\cdot)$ and $\Phi(\cdot)$ the probability density function and cumulative distribution function of the standard normal distribution, respectively.
@@ 2.2 Ideal Estimands
For each A/B test $i$, the expected performance of a methodology $\kappa$ under performance measure $p$ given a true impact of $\Delta_{i}=\delta_{i}$ is given by:
@@
Here $\kappa(\hat{X}_{i})$ is the methodology’s output (estimate or launch decision) given data $\hat{X}_{i}$. The function $p(\kappa(\hat{X}_{i}),\delta_{i})$ measures how well this output performs given a true impact $\delta_{i}$. Taking the expectation over the sampling variability induced by $F$ averages out the randomness in the data $\hat{X}_{i}$, yielding the expected performance $Y_{i}(\kappa)$ for a given true impact.
While $Y_{i}(\kappa)$ captures the expected performance of $\kappa$ for a given test, it varies across tests due to differences in the true impacts $\Delta_{i}$:
@@
Example 3 (Illustrating performance estimands for estimators): Assume that $\tau_{i}=\tau$ for all i=1,,Ii=1,...,I and that $G\sim N(0,\sigma^{2})$. Consider evaluating the relative performance of two estimators under squared error $p(\kappa(\hat{x}_{i}),\delta_{i})=(\kappa(\hat{x}_{i})-\delta_{i})^{2}$: the unbiased estimator $\kappa^{1}(\hat{x}_{i})=\hat{\delta}_{i}$ and the Bayes estimator $\kappa^{2}(\hat{x}_{i})=\sigma^{2}(\sigma^{2}+\tau^{2})^{-1}\hat{\delta}_{i}$. In this case, we have that:
@@
consistent with the fact that the Bayes estimator minimizes mean squared error ([casella_berger_2002]).
Example 4 (Illustrating performance estimands for decision rules): Assume that $\tau_{i}=\tau$ for all i=1,,Ii=1,...,I and that $G\sim N(0,\sigma^{2})$. Consider evaluating the relative performance of two decision rules under launch-only decision value $p(\kappa(\hat{x}_{i}),\delta_{i})=\kappa(\hat{x}_{i})\cdot\delta_{i}$: a frequentist decision rule $\kappa^{1}(\hat{x}_{i})=1(\hat{\delta}_{i}>\tau\cdot\Phi^{-1}(0.975))$ which launches treatments when the estimated impact is positive and statistically significant at the 5 percent level and a Bayes decision rule $\kappa^{2}(\hat{x}_{i})=1(\sigma^{2}(\sigma^{2}+\tau^{2})^{-1}\hat{\delta}_{i}>0)$ which launches treatments when the posterior mean is positive. In this case, we have that:
@@
consistent with the fact that the Bayes decision rule maximizes mean launch-only decision value ([azevedo_etal_2020]).
Note that these examples assume that the distribution of true impacts $G$ is normal only to help make the estimands concrete. Our sample splitting framework, introduced below, is designed to evaluate performance of methodologies without requiring any assumptions about $G$.
@@ 2.3 Sample Splitting
We evaluate the performance of methodologies via sample splitting, where we partition each A/B test into two splits: one to implement the proposed methodology, the other to construct an independent, unbiased proxy for the true impact. This approach enables performance measurement when true impacts are unobserved ([tripuraneni_etal_2023]). However, because the proposed methodology is implemented on only a split of the sample, we cannot directly estimate the ideal estimands in section 2.2. In what follows, we outline a general approach to generating sample-split data, which applies to any methodology. We then show that this approach can be simplified for “plug-in methodologies” that depend on the data only through a set of sufficient statistics. With this data foundation, we then define sample-split analogues of the ideal estimands, introduce corresponding unbiased estimators, and outline an approach to statistical inference.
@@ 2.3.1 Data Generation: General Approach
Because the methodology $\kappa(\hat{X}_{ias})$ and the proxy for the true impact $\hat{\Delta}_{ibs}$ are based on disjoint samples of units, they are statistically independent conditional on the true impact $\Delta_{i}=\delta_{i}$:
$\kappa\left(\hat{X}_{ias}\right)\perp\hat{\Delta}_{ibs}\>\big|\>\Delta_{i}=\delta_{i}$ (7)
@@
$\hat{\Delta}_{ibs}|\Delta_{i}=\delta_{i},\tau_{i}^{2}\overset{\text{ind}}{\sim}N\left(\delta_{i},\tau_{ibs}^{2}\right)$ (8)
@@ 2.3.2 Data Generation: Plug-in Methodologies
Proposition 1 is similar in spirit to [lewis_2025], who generate “doppelganger” observations without access to unit-level data by deconvolving the overall estimates for a test.
@@ 2.3.3 Estimands
To evaluate performance of methodologies under sample splitting, we introduce generalized, sample-splitting analogues of the ideal estimands introduced in 2.2. These estimands reflect that performance depends not only on the methodology $\kappa$ and performance function $p$, but also on the share of the sample $\alpha$ used in implementing the methodology. Given $\alpha$, the expected performance of a methodology $\kappa$ under performance measure $p$ given a true impact of $\Delta_{i}=\delta_{i}$ is given by:
@@
Just like its ideal counterpart $Y_{i}(\kappa)$, $Y_{i}(\kappa,\alpha)$ varies across tests due to differences in true impacts $\Delta_{i}$:
@@ 2.3.4 Estimators and Inference
The variance $\Omega_{i}^{2}\left(\kappa,\alpha,S\right)$ reflects both sampling variability in the full-sample data $\hat{X}_{i}$ given the true impact $\Delta_{i}=\delta_{i}$ and variability in the sample-split data $(\hat{X}_{ias},\hat{X}_{ibs})$ given the full-sample data.
@@
Decision rule Decision value $E_{F}(\kappa(\hat{X}_{ias})\cdot\delta_{i}-(1-\kappa(\widehat{X}_{ias}))\cdot\delta_{i})$ $\kappa(\hat{X}_{ias})\cdot\hat{\delta}_{ibs}-(1-\kappa(\hat{X}_{ias}))\cdot\hat{\delta}_{ibs}$
@@ 4 Takeaways for Practitioners
We conclude by distilling our theoretical and simulation findings into four concrete recommendations for sample-splitting practitioners:
@@ Appendix A Appendix: Proof of Claims
Proof: Assume that equations (7) and (8) hold. For bias, the claim then follows immediately from the unbiasedness of $\hat{\Delta}_{ibs}$. For squared error, we follow [tripuraneni_etal_2023], noting that:
@@
where equation (34) follows from the independence of $\kappa(\hat{X}_{ias})$ and $\hat{\Delta}_{ibs}$ conditional on $\Delta_{i}=\delta_{i}$ and the unbiasedness of $\hat{\Delta}_{ibs}$. The claim then follows in equation (35) from the definition of the sampling variance of $\hat{\Delta}_{ibs}$. For launch-only decision value, we follow [tripuraneni_etal_2023], noting that:
@@
$=E_{F}\left(\kappa\left(\hat{X}_{ias}\cdot\delta_{i}\right)\right)$ (37)
@@
$\widehat{\zeta}^{2}(\kappa,\alpha,S)=\frac{1}{I}\sum_{i=1}^{I}(\hat{Y}_{i}(\kappa,\alpha,S)-\hat{\theta}(\kappa,\alpha,S))^{2}$ (51)

### v2

@@ 1 Introduction
Measuring the impact of proposed improvements to the way in which A/B tests are analyzed and used to inform decision making is critical for large-scale A/B testing products. Over the last few decades, there have been several proposed improvements to A/B testing methodologies, including new estimators of impacts (for example, [9], [10], and [8]) and new rules for deciding when to launch (for example, [11], [2] and [3]). The scope of proposed improvements continues to expand, with more recent work exploring AI-driven approaches (for example, [7] and [6]). A/B testing products must determine which of these methodologies deliver meaningful value relative to their implementation costs.
@@
Sample splitting offers a potential solution to this problem ([14]), but its theoretical foundation remains undeveloped. Under sample splitting, each A/B test is divided into two, non-overlapping splits: one to implement the proposed methodology, the other to construct an independent, unbiased proxy for the true impact. Comparing these two quantities across A/B tests allows us to measure the average performance of the proposed methodology, without making strong distributional assumptions ([14]). However, the statistical properties of these sample-split estimators have yet to be characterized formally: their bias, variance, and asymptotic distributions remain unknown, and there are no established methods for conducting statistical inference. Without such foundations, products lack guidance on exactly how to use sample splitting, limiting its adoption in practice.
@@
More generally, our work contributes to the literature on cross-validation and resampling methods ([13]). Like cross-validation, sample splitting separates the application of a methodology from its evaluation by partitioning data. The key difference is that in cross-validation the prediction target is observed directly, whereas in our setting the prediction target is itself an estimate subject to sampling variation. To our knowledge, [14] were the first to apply the idea of sample splitting in the A/B testing context, building upon earlier work in the estimation of heterogeneous treatment effects by [1] and [5]. [14] demonstrate the potential of sample splitting in evaluating the performance of A/B testing methodologies but leave its statistical properties largely unformalized. [^1] The authors acknowledge that their work is guided primarily by intuition and heuristics rather than principled theory. For example, they note the use “$t$-test heuristics” (page 7) and suggest that the confidence intervals they report “should be interpreted cautiously” (page 11). We extend their work by introducing a theoretical framework for sample splitting in A/B testing environments, characterizing both what these methods can measure and how to conduct valid inference.
@@ 2.1 Context
We consider a set of $i\in\{1,...,I\}$ feature changes, each evaluated through an A/B test where the change is exposed to a random subset of units. Each feature change has a true impact $\Delta_{i}$, which is a random variable with realized values $\delta_{i}$. Each test is further characterized by latent variables $Z_{i}$, with realized values $z_{i}$, capturing features of the testing environment such as sample size and outcome variability. The pairs $(\Delta_{i},Z_{i})$ are independently and identically distributed according to an unknown distribution $G$.
Conditional on the true impact $\Delta_{i}=\delta_{i}$ and test environment $Z_{i}=z_{i}$, the A/B test generates random data $\hat{X}_{i}$ according to a sampling distribution $F_{i}$, with realized values $\hat{x}_{i}$. To simplify notation, we suppress the dependence on $i$ and write $F$ in place of $F_{i}$. This sampling distribution reflects the sampling of units and random assignment of treatment in each test and may vary across tests through the true impact and test environment $(\Delta_{i},Z_{i})$.
We define an A/B testing methodology $\kappa(\hat{x}_{i})$ to be a mapping of the data $\hat{x}_{i}$ to a real-valued output. The methodology could be an estimator of the true impact $\delta_{i}$ or a decision rule that determines whether to launch the given treatment ($\kappa(\hat{x}_{i})=1$) or not ($\kappa(\hat{x}_{i})=0$).
Given the randomization of the feature change and the central limit theorem, we assume that each test yields an unbiased estimator $\hat{\Delta}_{i}$ of the impact $\delta_{i}$ that is normally distributed, with realized values $\hat{\delta}_{i}$ and sampling variance $\tau_{i}^{2}=\tau_{i}^{2}(\delta_{i},z_{i})$:
$\hat{\Delta}_{i}|\Delta_{i}=\delta_{i},Z_{i}=z_{i}\overset{\text{ind}}{\sim}N(\delta_{i},\tau_{i}^{2})$ (1)
Following standard practice in the A/B testing literature (see, for example, [3]), we treat $\tau_{i}^{2}$ as known for each test, while allowing it to vary across tests.
@@
In what follows, for any random variable $K_{i}$, we write $F_{K}$ to denote its distribution induced by the sampling distribution $F$ and $G_{K}$ to denote its distribution induced by the distribution $G$. We write $E_{F}(\cdot)$ and $\text{Var}_{F}(\cdot)$ for the expectation and variance with respect to randomness induced by $F$, and $E_{G}(\cdot)$ and $\text{Var}_{G}(\cdot)$ for the expectation and variance with respect to randomness induced by $G$. Finally, we write $E(\cdot)$ and $\text{Var}(\cdot)$ for the expectation and variance with respect to both sources of randomness, integrating over both $F$ and $G$. We denote by $\phi(\cdot)$ and $\Phi(\cdot)$ the probability density function and cumulative distribution function of the standard normal distribution, respectively.
@@ 2.2 Ideal Estimands
For each A/B test $i$, the expected performance of a methodology $\kappa$ under performance measure $p$ given a true impact of $\Delta_{i}=\delta_{i}$ and test environment $Z_{i}=z_{i}$ is given by:
@@
Here $\kappa(\hat{X}_{i})$ is the methodology’s output (estimate or launch decision) given data $\hat{X}_{i}$. The function $p(\kappa(\hat{X}_{i}),\delta_{i})$ measures how well this output performs given a true impact $\delta_{i}$. Taking the expectation over the sampling variability induced by $F$ averages out the randomness in the data $\hat{X}_{i}$, yielding the expected performance $Y_{i}(\kappa)$ for a given true impact and test environment.
While $Y_{i}(\kappa)$ captures the expected performance of $\kappa$ for a given test, it varies across tests due to variation in $(\Delta_{i},Z_{i})$:
@@
Example 3 (Illustrating performance estimands for estimators): Assume that $G$ is such that $\tau_{i}=\tau$ for all i=1,,Ii=1,...,I and $\Delta_{i}\sim N(0,\sigma^{2})$. Consider evaluating the relative performance of two estimators under squared error $p(\kappa(\hat{x}_{i}),\delta_{i})=(\kappa(\hat{x}_{i})-\delta_{i})^{2}$: the unbiased estimator $\kappa^{1}(\hat{x}_{i})=\hat{\delta}_{i}$ and the Bayes estimator $\kappa^{2}(\hat{x}_{i})=\sigma^{2}(\sigma^{2}+\tau^{2})^{-1}\hat{\delta}_{i}$. In this case, we have that:
@@
consistent with the fact that the Bayes estimator minimizes mean squared error ([4]).
Example 4 (Illustrating performance estimands for decision rules): Assume that $G$ is such that $\tau_{i}=\tau$ for all i=1,,Ii=1,...,I and $\Delta_{i}\sim N(0,\sigma^{2})$. Consider evaluating the relative performance of two decision rules under launch-only decision value $p(\kappa(\hat{x}_{i}),\delta_{i})=\kappa(\hat{x}_{i})\cdot\delta_{i}$: a frequentist decision rule $\kappa^{1}(\hat{x}_{i})=1(\hat{\delta}_{i}>\tau\cdot\Phi^{-1}(0.975))$ which launches treatments when the estimated impact is positive and statistically significant at the 5 percent level and a Bayes decision rule $\kappa^{2}(\hat{x}_{i})=1(\sigma^{2}(\sigma^{2}+\tau^{2})^{-1}\hat{\delta}_{i}>0)$ which launches treatments when the posterior mean is positive. In this case, we have that:
@@
consistent with the fact that the Bayes decision rule maximizes mean launch-only decision value ([3]).
Note that these examples make restrictive assumptions about the distribution $G$ only to help make the estimands concrete. Our sample splitting framework, introduced below, is designed to evaluate performance of methodologies without requiring any assumptions about $G$.
@@ 2.3 Sample Splitting
We evaluate the performance of methodologies via sample splitting, where we partition each A/B test into two splits: one to implement the proposed methodology, the other to construct an independent, unbiased proxy for the true impact. This approach enables performance measurement when true impacts are unobserved ([14]). However, because the proposed methodology is implemented on only a split of the sample, we cannot directly estimate the ideal estimands in section 2.2. In what follows, we outline a general approach to generating sample-split data, which applies to any methodology. We then show that this approach can be simplified for “plug-in methodologies” that depend on the data only through a set of sufficient statistics. With this data foundation, we then define sample-split analogues of the ideal estimands, introduce corresponding unbiased estimators, and outline an approach to statistical inference.
@@ 2.3.1 Data Generation: General Approach
Because the methodology $\kappa(\hat{X}_{ias})$ and the proxy for the true impact $\hat{\Delta}_{ibs}$ are based on disjoint samples of units, they are statistically independent conditional on the true impact $\Delta_{i}=\delta_{i}$ and testing environment $Z_{i}=z_{i}$:
$\kappa\left(\hat{X}_{ias}\right)\perp\hat{\Delta}_{ibs}\>\big|\>\Delta_{i}=\delta_{i},Z_{i}=z_{i}$ (7)
@@
$\hat{\Delta}_{ibs}|\Delta_{i}=\delta_{i},Z_{i}=z_{i}\overset{\text{ind}}{\sim}N\left(\delta_{i},\tau_{ibs}^{2}\right)$ (8)
@@ 2.3.2 Data Generation: Plug-in Methodologies
Proposition 1 is similar in spirit to [12], who generate “doppelganger” observations without access to unit-level data by deconvolving the overall estimates for a test.
@@ 2.3.3 Estimands
To evaluate performance of methodologies under sample splitting, we introduce generalized, sample-splitting analogues of the ideal estimands introduced in section 2.2. These estimands reflect that performance depends not only on the methodology $\kappa$ and performance function $p$, but also on the share of the sample $\alpha$ used in implementing the methodology. Given $\alpha$, the expected performance of a methodology $\kappa$ under performance measure $p$ given a true impact of $\Delta_{i}=\delta_{i}$ and testing environment $Z_{i}=z_{i}$ is given by:
@@
Just like its ideal counterpart $Y_{i}(\kappa)$, $Y_{i}(\kappa,\alpha)$ varies across tests due to variation in $(\Delta_{i},Z_{i})$:
@@ 2.3.4 Estimators and Inference
The variance $\Omega_{i}^{2}\left(\kappa,\alpha,S\right)$ reflects both sampling variability in the full-sample data $\hat{X}_{i}$ given the true impact $\Delta_{i}=\delta_{i}$ and testing environment $Z_{i}=z_{i}$ and variability in the sample-split data $(\hat{X}_{ias},\hat{X}_{ibs})$ given the full-sample data.
@@
Decision rule Decision value $E_{F}(\kappa(\hat{X}_{ias})\cdot\delta_{i}-(1-\kappa(\hat{X}_{ias}))\cdot\delta_{i})$ $\kappa(\hat{x}_{ias})\cdot\hat{\delta}_{ibs}-(1-\kappa(\hat{x}_{ias}))\cdot\hat{\delta}_{ibs}$
@@ 4 Takeaways for Practitioners
We conclude by distilling our theoretical and simulation findings into five recommendations for sample-splitting practitioners:
@@ Appendix A Appendix: Proof of Claims
Proof: Assume that equations (7) and (8) hold. For bias, the claim then follows immediately from the unbiasedness of $\hat{\Delta}_{ibs}$. For squared error, we follow [14], noting that:
@@
where equation (34) follows from the independence of $\kappa(\hat{X}_{ias})$ and $\hat{\Delta}_{ibs}$ conditional on $(\delta_{i},z_{i})$ and the unbiasedness of $\hat{\Delta}_{ibs}$. The claim then follows in equation (35) from the definition of the sampling variance of $\hat{\Delta}_{ibs}$. For launch-only decision value, we follow [14], noting that:
@@
$=E_{F}\left(\kappa\left(\hat{X}_{ias}\right)\cdot\delta_{i}\right)$ (37)